# bibliography.bib

### v2

@@ 3 Finite element methods for PDEs with random data
As a use case for inexact computations within multilevel sampling methods, we now consider the multilevel Monte Carlo method and a standard elliptic PDE with random data discretised via finite elements (FE). Problems of this kind arise, for example, in geosciences, namely, in the study of groundwater flow; see [scheidegger1957physics, hoeksema1985analysis, freeze1975stochastic], and the references therein.
@@ 3.1 An elliptic PDE with random data and its FE solution
Under the assumption (4) the unique solvability of the AVP with random data (Problem 3.1) can be proved sample-wise using the Lax-Milgram lemma in the standard way; see [babuska2004galerkin]. The coercivity and continuity constants in the Lax-Milgram lemma do not depend on $\omega$. More details regarding the analysis of Problem 3.1 in the stochastic context can be found in [babuska2004galerkin, barth2011multi]. Let us note that if the bounds in (4) are weakened and depend on $\omega$, the analysis is still possible but it becomes more complicated; see [teckentrup2013further] for a thorough discussion.
@@
In order to solve Problem 3.1, we use conforming FEs on a shape-regular and quasi-uniform family of triangulations of the domain $D\subset\mathbb{R}^{n}$ (in our examples $n=2$) with a mesh parameter $h>0$. We employ (piecewise linear) $\mathcal{P}_{1}$ Lagrange elements to compute the discrete solution $u_{h}(\cdot,\omega)$ to the AVP with random data (Problem 3.1) in the FE space $V_{h}\subset V$. Using a basis of $V_{h}$ consisting of hat functions $\phi_{j}$, this is equivalent to solving a linear system $A(\omega)x(\omega)=b(\omega)$ with a positive definite matrix $A(\omega)$; see [ern2004theory] for details.
@@ 3.2 Approximate FE solutions using inexact solvers
For $\omega\in\Omega$ fixed, the bound follows from [ern2004theory, Proposition 9.19], with a constant independent of $u$ and $h$. To see that $C$ is also independent of $\omega$, we write out the constant $C$ from [ern2004theory, Proposition 9.19] explicitly:
@@
Here, $c_{P}$ is the Poincaré constant, $\alpha$ is the coercivity constant of the bilinear form $\mathcal{A}$ and $\kappa(\mathcal{M}_{t})$ is the condition number of the mass matrix $\mathcal{M}_{t}$. It follows from the definition of $\mathcal{M}_{t}$ that $\kappa(\mathcal{M}_{t})$ is independent of $\omega$; see [ern2004theory, Section 9.1.3]. Due to (4), the coercivity constant $\alpha$ is also independent of $\omega$. This completes the proof. ∎
Thus, to control the error $\|u_{h}(\cdot,\omega)-\widehat{u}_{h}(\cdot,\omega)\|_{H_{0}^{1}(D)}$ it suffices to control the relative residual $\|{r(\omega)}\|_{2}/\|{b(\omega)}\|_{2}$ of the resulting linear system. In the following we derive the convergence rate of iterative refinement from Section 2.3 applied to the FE system from Problem 3.1. This will guide our choice of the precisions in iterative refinement (Algorithm 2.1). To proceed with our analysis, we need some auxiliary inequalities from [ern2004theory], which we summarize here.
@@
where $\kappa_{2}$ is the condition number of $A$ with respect to the spectral norm and the generic constant $c$ is independent of the discretisation parameter $h$ and of $\omega$.
@@
The first claim follows from the proof of [ern2004theory, Theorem 9.11] (the last inequality in the first part) together with [ern2004theory, Theorem 9.8]. The inequality from [ern2004theory, Theorem 9.11] gives us a bound on $\|{A}\|_{2}$ using $h$ and the eigenvalues of the mass matrix, which then can be bounded using [ern2004theory, Theorem 9.8].
The second claim follows from [ern2004theory, Theorem 9.11] (with $s=t=1$) as well as [ern2004theory, Example 9.13]. The $\omega$-independence of the constant $c$ follows as in Lemma 3.2 from the definition of the stiffness matrix $A$ and from (4). ∎
As a consequence of Lemma 3.3 we formulate the following corollary about convergence properties of iterative refinement for the FE systems. It follows immediately from [carson2018accelerating, Corollary 4.2] using the inequalities from Lemma 3.3 and the fact that $\|x(\omega)\|_{2}\leq\|A^{-1}(\omega)\|_{2}\|b(\omega)\|_{2}$. For the purpose of this corollary we make the reasonable assumption that $\|{\widehat{x}_{i}(\omega)}\|_{2}\approx\|{x(\omega)}\|_{2}$.
@@
Let $\phi$ be the approximate convergence rate of iterative refinement applied to the FE system $A(\omega)x(\omega)=b(\omega)$ corresponding to Problem 3.1. The solution eventually computed by Algorithm 2.1 is denoted by $\widehat{x}_{i}(\omega)$. Then
@@
    the factor $\phi$ satisfies $\phi\leq c\varepsilon_{s}(1+h^{-2})$ and
@@
    $\frac{\|{b(\omega)-A(\omega)\widehat{x}_{i}(\omega)}\|_{2}}{\|{b(\omega)}\|_{2}}\leq cp_{z}\varepsilon(1+h^{-2})$,
with generic constants derived from [carson2018accelerating, Corollary 4.2] and Lemma 3.3.
This corollary allows us to analyse convergence properties of iterative refinement for our model problem. In particular, for a given value of the tolerance $\delta_{e}$, we are able to derive upper bounds on the precisions $\varepsilon_{r}$$\varepsilon$, and $\varepsilon_{f}$ in Algorithm 2.1 (recall that in our case $\varepsilon_{s}=\varepsilon_{f}$). The convergence rate $\phi$ has to be “sufficiently smaller than $1$” with the exact meaning of this statement given in [carson2018accelerating, Theorem 4.1]. The first bound in Corollary 3.4, thus leads to an upper bound for the factorisation precision $\varepsilon_{f}$. Using the second bound in Corollary 3.4, it is sufficient to set $cp\varepsilon(1+h^{-2})<\delta_{e}$, which yields a bound on the working precision $\varepsilon$. The residual precision only needs to satisfy the condition $\varepsilon_{r}\leq\varepsilon$, so that we can set $\varepsilon_{r}=\varepsilon$. Under these constraints on $\varepsilon_{r}$, $\varepsilon$, and $\varepsilon_{f}$, [carson2018accelerating, Corollary 4.2] guarantees that Algorithm 2.1 converges and Corollary 3.4 provides a bound on the convergence rate $\phi$.
@@ 4.1 Standard multilevel Monte Carlo
Multilevel Monte Carlo (MLMC) is one of the standard multilevel sampling methods to compute expectations for uncertainty quantification (UQ) in PDE applications, and a suitable example to show how inexact computations can be leveraged in that context. The main idea of MLMC is to optimally balance the sampling and discretisation errors of a hierarchy of approximate models; see [giles2015multilevel] for an overview. We formulate it here in full generality:
@@
In practice, we use an adaptive algorithm to determine the values $L$ and $\{N_{l}\}_{l=0}^{L}$ for a given tolerance TOL; see [cliffe2011multilevel, Section 2]). The adaptive MLMC algorithm aims to compute the optimal values of $L$ and $\{N_{l}\}_{l=0}^{L}$ by minimizing the computational cost for a given variance. To this end, the algorithm uses sample averages $\widehat{Y}_{l}$ (see (8)) and variance estimators
@@
It can be shown that three complexity regimes can be distinguished based on the values of $\beta$ and $\gamma$. The optimal performance of the MLMC algorithm is achieved when $\beta>\gamma$, i.e., when variance decays faster than cost increases. The full complexity analysis can be found in [cliffe2011multilevel, Theorem 1].
@@ 4.2 Computational error in multilevel Monte Carlo
In this section we present a convergence analysis taking into account the computational error in the MLMC method discussed in Section 4.1. Our analysis shows that especially on the coarser levels, it is sufficient to compute with relatively low effective precision. It suggests that significant gains in terms of both memory and computational time can be obtained depending on the model considered and its implementation. To this end, we propose a novel adaptive MLMC algorithm, which determines the minimum required numerical precision with no additional cost. This algorithm will be referred to as mixed precision multilevel Monte Carlo (MPML) for simplicity. The efficiency of the adaptive algorithm is demonstrated on numerous experiments in Section 5.
Let us start by defining the setting. As in Section 4.1, we assume $Q:\Omega\rightarrow\mathbb{R}$ to be a random variable and the goal is to compute its mean $\mathbb{E}[Q]$. This time we assume that we cannot evaluate $Q_{l}(\omega)$ exactly, but only an approximation $\tilde{Q}_{l}$ of $Q_{l}$. This might be due to finite precision arithmetic for the model evaluation or for generating the random variable. Using $\tilde{Q_{l}}$ instead of $Q_{l}$ in the MLMC estimator (8) gives us what will be referred to as the mixed precision multilevel Monte Carlo estimator and will be denoted by $\widehat{Q}_{L,\{N_{l}\}}^{\text{MPML}}$. The crucial difference between the MPML estimator and the standard MLMC estimator is the error model which is used. For MPML we propose an additive error model stated below. Using the bias-variance decomposition we can also quantify the overall error of the MPML estimator.
@@
Let $m\in\mathbb{N}$, $m>1$ and assume that $\delta_{0},\delta_{1},\ldots$ is a sequence of parameters characterising computational error with $01>\delta_{l}>$ 0. Assume that there exist $\alpha_{1},\beta_{1},\alpha_{2},\beta_{2}>0$ such that
@@
The inequality follows straightforwardly from the bias-variance decomposition (see [cliffe2011multilevel]) and the assumptions (11) and (12). ∎
This general error estimate can be used in practice to compute a bound on the computational error so that the overall MSE does not exceed a certain tolerance. A concrete example of what form the parameter $\delta_{l}$ can take will be discussed in the next section. For the purpose of this general error estimate we have not assumed that the computational error decays. An example of this would be the situation when we have a hierarchy of approximations of a linear PDE and solve the resulting system of linear algebraic equations on each level using an iterative solver with a fixed number of iterations, or a sparse direct solver in finite precision. Then the error of the model decays with increasing $l$, but we expect the computational error to grow. This is in accordance with what has been observed in [giles2024rounding, Figure 3.1].
4.3 Application to the elliptic PDE problem
Problem 4.2.
Under assumptions (4) and (5), it can be shown that when MLMC is applied to Problem 4.2, we obtain (10) with $\alpha=2$ and $\beta=4$. Generally $\gamma$ depends on the linear solver used and we discuss it in our numerical experiments. If an optimal linear solver is used (e.g., multigrid), one has $\gamma=2$; see [cliffe2011multilevel].
In our PDE problem, we assume a hierarchy of discrete models based on a uniform refinement with a factor $m>1$ of an inital mesh with mesh size $h_{0}$. Assumptions (11) and (12) may then be rewritten as
@@ 4.3 Application to the elliptic PDE problem
The values of the constants $\alpha_{1}$, $\alpha_{2}$, $\beta_{1}$, and $\beta_{2}$ are given by the following lemma.
Lemma 4.3.
$|\mathbb{E}[\tilde{Q}_{l}-Q]|$ $=O(h_{l}^{2}+\delta_{l}),$ (15)
$\operatorname{\textrm{var}}[\tilde{Y}_{l}]$ $=O(h_{l}^{4}+\delta_{l}^{2}).$ (16)
Proof.
Using Jensen’s inequality the bias error in (15) can be decomposed as
From [cliffe2011multilevel, Section 3] it follows that
$\|G(u)-G(u_{h})\|_{L^{1}(\Omega,\mathbb{R})}\leq Ch^{2},$ (18)
The variance bound in (16) can be shown similarly. Let us estimate
$\leq\mathbb{E}\bigl{[}\bigl{(}\tilde{Q}_{l}-Q_{h_{l}}+Q_{h_{l}}-Q+Q-Q_{h_{l-1}}+Q_{h_{l-1}}-\tilde{Q}_{l-1}\bigr{)}^{2}\bigr{]}.$
Using the same technique as in [cliffe2011multilevel], we obtain an estimate of the form
$\operatorname{\textrm{var}}[\tilde{Y}_{l}]\leq$ $C\bigl{(}\mathbb{E}[(\tilde{Q}_{l}-Q_{h_{l}})^{2}]+\mathbb{E}[(Q_{h_{l}}-Q)^{2}]$
$+\mathbb{E}[(Q-Q_{h_{l-1}})^{2}]+\mathbb{E}[(Q_{h_{l-1}}-\tilde{Q}_{l-1})^{2}]\bigr{)}.$ (20)
As in [cliffe2011multilevel], the quantity $\mathbb{E}[(Q_{h_{l}}-Q)^{2}]+\mathbb{E}[(Q-Q_{h_{l-1}})^{2}]$ can be estimated by
$\mathbb{E}[(Q_{h_{l}}-Q)^{2}]+\mathbb{E}[(Q-Q_{h_{l-1}})^{2}]\leq Ch_{l}^{4}.$ (21)
This lemma allows us to formulate a specific version of Theorem 4.1 regarding the error of the MPML estimator applied to Problem 4.2 discretised using finite elements. We summarize this in the following corollary.
Corollary 4.4 (Error of the MPML FEM).
Proof.
The claim follows from Lemma 4.3 and Theorem 4.1. ∎
In this section we have not focused on the cost of the MPML method. This is because the asymptotic cost remains the same as for the standard MLMC method as long as $\delta_{L}=O(h_{L}^{\alpha})$ and $\delta_{l}^{2}=O(h_{l}^{\beta})$. However, due to the use of inexact solvers, significant time and memory gains can be obtained.

### v3

@@ 3 Finite element methods for PDEs with random data
As a use case for inexact computations within multilevel sampling methods, we now consider the multilevel Monte Carlo method and a standard elliptic PDE with random data, discretised via finite elements (FE). Problems of this kind arise, for example, in geosciences, namely, in the study of groundwater flow; see [28, 23, 13], and the references therein.
@@ 3.1 An elliptic PDE with random data and its FE solution
Under the assumption (4) the unique solvability of the AVP with random data (Problem 3.1) can be proved sample-wise using the Lax-Milgram lemma in the standard way; see [1]. The coercivity and continuity constants in the Lax-Milgram lemma do not depend on $\omega$. More details regarding the analysis of Problem 3.1 in the stochastic context can be found in [1, 3]. Let us note that if the bounds in (4) are weakened and depend on $\omega$, the analysis is still possible but it becomes more complicated; see [30] for a thorough discussion.
@@
In order to solve Problem 3.1, we use conforming FEs on a shape-regular and quasi-uniform family of triangulations of the domain $D\subset\mathbb{R}^{n}$ (in our examples $n=2$) with a mesh parameter $h>0$. We employ (piecewise linear) $\mathcal{P}_{1}$ Lagrange elements to compute the discrete solution $u_{h}(\cdot,\omega)$ to the AVP with random data (Problem 3.1) in the FE space $V_{h}\subset V$. Using a basis of $V_{h}$ consisting of hat functions $\phi_{j}$, this is equivalent to solving a linear system $A(\omega)x(\omega)=b(\omega)$ with a positive definite matrix $A(\omega)$; see [12] for details.
@@ 3.2 Approximate FE solutions using inexact solvers
For $\omega\in\Omega$ fixed, the bound follows from [12, Proposition 9.19], with a constant independent of $u$ and $h$. To see that $C$ is also independent of $\omega$, we write out the constant $C$ from [12, Proposition 9.19] explicitly:
@@
Here, $c_{P}$ is the Poincaré constant, $\alpha$ is the coercivity constant of the bilinear form $\mathcal{A}$ and $\kappa(\mathcal{M}_{t})$ is the condition number of the mass matrix $\mathcal{M}_{t}$. It follows from the definition of $\mathcal{M}_{t}$ that $\kappa(\mathcal{M}_{t})$ is independent of $\omega$; see [12, Section 9.1.3]. Due to (4), $\alpha$ is also independent of $\omega$. This completes the proof. ∎
Thus, to control the error $\|u_{h}(\cdot,\omega)-\widehat{u}_{h}(\cdot,\omega)\|_{H_{0}^{1}(D)}$ it suffices to control the relative residual $\|{r(\omega)}\|_{2}/\|{b(\omega)}\|_{2}$ of the resulting linear system. In the following we derive the convergence rate of iterative refinement from Section 2.3 applied to the FE system from Problem 3.1. This will guide our choice of the precisions in iterative refinement (Algorithm 2.1). To proceed with our analysis, we need some auxiliary inequalities from [12], which we summarize here.
@@
where $\kappa_{2}$ is the condition number of AA with respect to the spectral norm and the generic constant $c$ is independent of the discretisation parameter $h$ and of $\omega$.
@@
The first claim follows from the proof of [12, Theorem 9.11] (the last inequality in the first part) together with [12, Theorem 9.8]. The inequality from [12, Theorem 9.11] gives us a bound on $\|{A}\|_{2}$ using $h$ and the eigenvalues of the mass matrix, which then can be bounded using [12, Theorem 9.8].
The second claim follows from [12, Theorem 9.11] (with $s=t=1$) as well as [12, Example 9.13]. The $\omega$-independence of the constant $c$ follows as in Lemma 3.2 from the definition of the stiffness matrix $A$ and from (4). ∎
As a consequence of Lemma 3.3 we formulate a corollary about the convergence of iterative refinement for the FE system. It follows immediately from [6, Corollary 4.2] using the inequalities from Lemma 3.3 and the fact that $\|x(\omega)\|_{2}\leq\|A^{-1}(\omega)\|_{2}\|b(\omega)\|_{2}$. For this corollary we make the reasonable assumption that the approximate solution $x_{i}(\omega)$ after $i$ steps of iterative refinement satisfies $\|{x_{i}(\omega)}\|_{2}\approx\|{x(\omega)}\|_{2}$. The analysis in [6] is in the maximum norm $\|\cdot\|_{\infty}$, but it is easily adapted to $\|\cdot\|_{2}$.
@@
Let $A(\omega)x(\omega)=b(\omega)$ be as in Lemma 3.3 and let $c_{1}$ and $c_{2}$ be as in [6, eq. (2.4)]. If $(c_{1}\kappa_{2}(A(\omega))+c_{2})\varepsilon_{s}\ll 1$, then there exists a constant $c>0$ such that the residual in iterative refinement (Algorithm 2.1) satisfies
$\|r_{i}(\omega)\|_{2}\leq c(1+h^{-2})\max\Big(\varepsilon_{s}\|r_{i-1}(\omega)\|_{2},\varepsilon\|{b(\omega)}\|_{2}\Big).$ (8)
where $c$ only depends on generic constants in [6, Corollary 4.2] and Lemma 3.3.
This corollary allows us to give a priori upper bounds on the precisions $\varepsilon_{r}$, $\varepsilon$, and $\varepsilon_{f}$ in Algorithm 2.1 that guarantee that the stopping criterion $\|{r_{i}(\omega)}\|_{2}/\|{b(\omega)}\|_{2}<\delta_{e}$ in Line 6 of Algorithm 2.1 is satisfied. In particular, it suffices that $\varepsilon_{r}$, $\varepsilon$, and $\varepsilon_{f}$ satisfy the following three conditions:
@@
    $(c_{1}\kappa_{2}(A(\omega))+c_{2})\varepsilon_{s}\ll 1$,
@@
    $c(1+h^{-2})\varepsilon<\delta_{e}$,
  3. 3.
    $\varepsilon_{r}\leq\varepsilon$.
Under these conditions, Algorithm 2.1 converges to the required tolerance $\delta_{e}$ and (8) provides a bound on the convergence rate.
@@ 4.1 Standard multilevel Monte Carlo
Multilevel Monte Carlo (MLMC) is one of the standard multilevel sampling methods to compute expectations for UQ in PDE applications, and a suitable example to show how inexact computations can be leveraged in that context. The main idea of MLMC is to optimally balance the sampling and discretisation errors of a hierarchy of approximate models; see [14] for an overview. We formulate it here in full generality:
@@
In practice, we use an adaptive algorithm to determine the values $L$ and $\{N_{l}\}_{l=0}^{L}$ for a given tolerance TOL; see [8, Section 2]). The adaptive MLMC algorithm aims to compute the optimal values of $L$ and $\{N_{l}\}_{l=0}^{L}$ by minimizing the computational cost for a given variance. To this end, the algorithm uses sample averages $\widehat{Y}_{l}$ (see (9)) and variance estimators
@@
It can be shown that three complexity regimes can be distinguished based on the values of $\beta$ and $\gamma$. The optimal performance of the MLMC algorithm is achieved when $\beta>\gamma$, i.e., when variance decays faster than cost increases. The full complexity analysis can be found in [8, Theorem 1].
@@ 4.2 Computational error in multilevel Monte Carlo
In this section we present a convergence analysis taking into account the computational error in the MLMC method discussed in Section 4.1. Our analysis shows that especially on the coarser levels, it is sufficient to compute with relatively low effective precision. Significant gains both in terms of memory and computational time can be obtained, depending on the model considered and its implementation. To this end, we propose a novel adaptive MLMC algorithm, which determines the minimum required numerical precision with no additional cost. This algorithm will be referred to as mixed-precision multilevel Monte Carlo (MPML) for simplicity. The efficiency of the adaptive algorithm is demonstrated on numerous experiments in Section 5.
Let us start by defining the setting. As in Section 4.1, we assume $Q:\Omega\rightarrow\mathbb{R}$ to be a random variable and the goal is to compute its mean $\mathbb{E}[Q]$. This time we assume that we cannot evaluate $Q_{l}(\omega)$ exactly, but only an approximation $\tilde{Q}_{l}$ of $Q_{l}$. This might be due to finite precision arithmetic for the model evaluation or for generating the random variable. Using $\tilde{Q_{l}}$ instead of $Q_{l}$ in the MLMC estimator (9) gives us what will be referred to as the mixed-precision multilevel Monte Carlo estimator and will be denoted by $\widehat{Q}_{L,\{N_{l}\}}^{\text{MPML}}$. The crucial difference between the MPML estimator and the standard MLMC estimator is the error model which is used. For MPML we propose an additive error model stated below. Using the bias-variance decomposition we can also quantify the overall error of the MPML estimator.
@@
Let $m\in\mathbb{N}$, $m>1$ and assume that $\delta_{0},\delta_{1},\ldots$ is a sequence of parameters characterising computational error with $1>\delta_{l}>0$. Assume that there exist $\alpha_{1},\beta_{1},\alpha_{2},\beta_{2}>0$ such that
@@
The inequality follows directly from the bias-variance decomposition
$\color[rgb]{0,0,0}\mathbb{E}\bigl[(\mathbb{E}[Q]-\widehat{Q}_{L,\{N_{l}\}}^{\text{MPML}})^{2}\bigr]=(\mathbb{E}[Q-\tilde{Q}_{L}])^{2}+\sum_{l=0}^{L}\frac{\operatorname{\textrm{var}}[\tilde{Y}_{l}]}{N_{l}}$}\
(see also [8]) and the assumptions (12) and (13). ∎
This general error estimate can be used in practice to compute a bound on the computational error so that the overall MSE does not exceed a certain tolerance. A concrete example of what form the parameter $\delta_{l}$ can take will be discussed in Section 4.5. For the purpose of this general error estimate we have not assumed that the computational error decays. An example of this would be the situation when we have a hierarchy of approximations of a linear PDE and solve the resulting system of linear algebraic equations on each level using an iterative solver with a fixed number of iterations, or a sparse direct solver in finite precision. Then the error of the model decays with increasing $l$, but we expect the computational error to grow. This is in accordance with what has been observed in [15, Figure 3.1].
In our PDE problem, we assume a hierarchy of discrete FE models based on uniform mesh refinement with a factor $m>1$ of an inital mesh with mesh size $h_{0}$, such that $h_{l}:=h_{0}m^{-l}$ for $l\in\mathbb{N}$. Assumptions 12 and 13 may then be rewritten as
@@
In this case Theorem 4.1 can be stated in the following form.
Corollary 4.2.
Under Assumption (14), the MSE of $\widehat{Q}_{L,\{N_{l}\}}^{\text{MPML}}$ satisfies
$\mathbb{E}\bigl[(\mathbb{E}[Q]-\widehat{Q}_{L,\{N_{l}\}}^{\text{MPML}})^{2}\bigr]\leq C\biggl(h_{L}^{2\alpha_{1}}+\sum_{l=0}^{L}\frac{h_{l}^{\beta_{1}}}{N_{l}}\Bigr)+\Bigl(h_{L}^{\alpha_{1}}\delta_{L}^{\alpha_{2}}+\delta_{L}^{2\alpha_{2}}+\sum_{l=0}^{L}\frac{\delta_{l}^{\beta_{2}}}{N_{l}}\Bigr)\biggr).$
Remark 4.3 (Asymptotic MPML cost).
Note that the asymptotic cost of the MPML method remains the same as for the standard MLMC method as long as $\delta_{L}^{\alpha_{2}}=O(h_{L}^{\alpha_{1}})$ and $\delta_{l}^{\beta_{2}}=O(h_{l}^{\beta_{1}})$. However, due to the use of lower precision arithmetic, the overall computational time can be reduced significantly which is the topic of the next section.